# A survey and benchmark of high-dimensional Bayesian optimization of discrete sequences

### v1

@@ 4.2 Benchmarking HDBO on PMO
| HillClimbing | Hvarfner’s VanillaBO | RandomLineBO | SAASBO | Turbo | Bounce | ProbRep | |
| Oracle () | |||||||
| albuterol_similarity | |||||||
| amlodipine_mpo | |||||||
| celecoxib_rediscovery | |||||||
| deco_hop | |||||||
| drd2_docking | |||||||
| fexofenadine_mpo | |||||||
| gsk3_beta | |||||||
| isomer_c7h8n2o2 | |||||||
| isomer_c9h10n2o2pf2cl | |||||||
| jnk3 | |||||||
| median_1 | |||||||
| median_2 | |||||||
| mestranol_similarity | |||||||
| osimetrinib_mpo | |||||||
| perindopril_mpo | |||||||
| ranolazine_mpo | |||||||
| rdkit_logp | |||||||
| rdkit_qed | |||||||
| sa_tdc | |||||||
| scaffold_hop | |||||||
| sitagliptin_mpo | |||||||
| thiothixene_rediscovery | |||||||
| troglitazone_rediscovery | |||||||
| valsartan_smarts | |||||||
| zaleplon_mpo | |||||||
@@
To benchmark the performance of HDBO on discrete sequences, we focus on the PMO benchmark (Gao et al.,, 2022; Huang et al.,, 2021; Brown et al.,, 2019). From the taxonomy (Section 3) we select frequently-tested methods from several families: Hvarfner’s VanillaBO, RandomLineBO, Turbo, BAxUS, SAASBO, Bounce, and ProbRep. We also include a HillClimbing baseline, which explores the input space by taking random Gaussian steps. All continuous solvers we start from the same initial data to ensure a fair comparison, and the discrete solvers (Bounce and ProbRep) initialize according to their implementations. We test the aforementioned methods on PMO (Gao et al.,, 2022; Huang et al.,, 2021), which requires a discrete representation of small molecules. Thus, we train two MLP VAEs on SELFIES representations of small molecules using Zinc250k (Irwin et al.,, 2020; Zhu et al.,, 2022). These generative models had 2 and 128 latent dimensions, allowing us to get an impression of how these models scale with dimensionality. We restrict sequences to be of length 70 (adding [nop] tokens for padding); the post-processing renders an alphabet of 64 SELFIES tokens. Details can be found in Sec. A.3.
The average best result over 3 runs (of maximum 300 iterations each) is presented in Tables 1 and 2 for 128D and 2D latent spaces respectively. We see a clear advantage in the optimizers that work on learned representations, instead of in discrete space. Such a discrepancy is to be expected: methods that optimize in latent space have been presented with information prior to their optimization campaigns, while methods like Bounce and ProbRep explore the whole discrete space. Further, the simple baseline is reliably beaten by the continuous alternatives in lower dimensions except Turbo, but this advantage is not as clear in the 128D case, signaling a more complex problem. Some of these tasks, however, are equally challenging for all solvers. deco_hop remains close to the original default value of 0.5, and there is no improvement over valsartan_smarts (which only REINVENT improves on in the original PMO results (Gao et al.,, 2022; Loeffler et al.,, 2024)). SAASBO did not scale gracefully with dimension in terms of training time; results are presented for the 2D case, and they are pending for 128D. This phenomenon has also been reported in comparisons made by Papenmeier et al., (2022) and Hvarfner et al., (2024). At the time of writing, these results are not comparable with PMO due to changes in scikit-learn’s loading of oracles. This issue has been raised in the TDC repository.444https://github.com/mims-harvard/TDC/issues/244
@@ 5 Conclusion
In this paper, we surveyed the field of high-dimensional Bayesian optimization (HDBO) focusing on discrete problems. This highlighted the need for (i) a novel taxonomy of the field that emphasizes the differences between methods that rely on unsupervised discrete information, and methods that optimize sequences directly, and (ii) a standardized framework for benchmarking these methods. We approach these in the form of two software tools: poli and poli-baselines. Using these tools, we implemented several HDBO methods and tested them in a standard benchmark for small molecule optimization. We find that optimizers using pre-trained latent-variable models have an edge over the other tested methods that work directly on sequence space. Our framework opens the door to fair and easily-replicable comparisons. We expect poli-baselines to be used by practitioners for running HDBO solvers in up-to-date environments compared across several tasks in our ongoing benchmark, which we plan to expand to other discrete objectives in poli.
@@
Although we taxonomize different families of HDBO methods, we have only benchmarked a subset; moreover, we only test a single setting: max. 300 iterations in PMO. This limits the generalizability of our conclusions. That being said, our benchmark is ongoing and we plan to include further experiments in the project’s website with, hopefully, participation from the community. Note that optimizing small molecules can open the door to both the potential of drug discovery, but also dual use (Urbina et al.,, 2022).
@@
The work was partly funded by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (NNF20OC0062606). RM is funded by the Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429) and VILLUM FONDEN (40516). SH was further supported by a research grant (42062) from VILLUM FONDEN as well as funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement 757360). WB was supported by VILLUM FONDEN (40578). This work was in part supported by the Pioneer Centre for AI (DRNF grant number P1). MGD thanks Sergio Garrido and Anshuk Uppal for feedback on early versions of this document, and Peter Mørch Groth for useful discussions.
@@ A.1 Methods Overview
| closed-form ASM | September 22, 2020 | Nathan Wycoff and Wild, (2021) | ✓ |
@@
| GaBO | November 22, 2021 | Jaquier et al., 2020b | ✓ |
@@ A.2 Reproducing results
We forked the official open source implementation of Bounce555https://github.com/LeoIV/bounce and added an interface between poli’s black boxes and their optimizer. Moreover, we made their implementation pip-installable. Bounce’s implementation is originally provided with an MIT license.
@@
provides an open-source implementation built on GPyTorch and BoTorch.666https://github.com/facebookresearch/bo_pr They provide a pip installable Python package which did not install until the dependencies mentioned above were fixed (to 1.11 and 0.7 respectively); further, the environment provided had to be updated by replacing the deprecated scikit-learn installation in a fork of their repository. After implementing an interface for poli black boxes, we relied on their script run_one_replication.py to implement a custom solver. The original PR code is provided with an MIT License.
@@
was implemented by following the tutorial on BoTorch,777https://github.com/pytorch/botorch/blob/main/tutorials/turbo_1.ipynb which can be found on GitHub under MIT License.
@@
is implemented using the Python package provided by the authors.888https://github.com/LeoIV/BAxUS Since their work builds on the original TuRBO code from Uber, it inherits their license.
@@ A.3 Training VAEs on SELFIES
We use PyTorch to implement a VAE with [one_hot_input, 1024, 512, 256, latent_dim] encoder and a symmetric decoder, using ReLU activations. The latent space prior used is a standard Gaussian, and the decoded distribution is a categorical. We use torch.distributions to compute ELBO losses without any weighting.999The exact implementation can be found here: https://github.com/MachineLearningLifeScience/hdbo_benchmark/blob/master/src/hdbo_benchmark/generative_models/vae_selfies.py
@@
We train for a maximum of 1000 epochs using early stopping with a patience of 50 epochs. Our batch sizes are , learning rates are , and optimizer is AdamW.101010The training script can be found here: https://github.com/MachineLearningLifeScience/hdbo_benchmark/blob/master/src/hdbo_benchmark/experiments/training_vae_on_zinc250k/train_vae.py
@@ A.4 Technical details on poli and poli-baselines
This section aims to give a short overview on the possibilities of poli and poli-baselines. For further information please refer to the full online documentation under https://machinelearninglifescience.github.io/poli-docs/.
@@
The aim of poli is to make it easier to benchmark new algorithms on new problems. A major issue on that endeavour are often conflicting package dependencies. When evaluating a brand-new algorithm on an older benchmark problem, it can simply be impossible to create a Python environment in which both can run. One may re-implement one or the other, though this is not only cumbersome but also error-prone. With the isolation mechanisms provided by poli this is no longer a problem. In particular, the isolation of the observer allows an easy comparison to additional algorithms or on other problems, without the need to change plotting scripts. poli provides a unified numpy (Harris et al.,, 2020) inferface, with inputs being numpy arrays of strings, and outputs being numpy arrays of floats. Finally, poli makes sure to log every call to the black box, as well as handling evaluation budgets—function calls are measured transparently and consistently across algorithms.
@@
The core object of poli is AbstractBlackBox, from which all black boxes inherit. The user implements a _black_box method that takes as input an array of strings x (as well as an optional context), and outputs the result of evaluating it. Listing 1 shows an example code snippet. At design time, a user registers the problem via a registry using the factory pattern (Gamma et al.,, 1995, p. 121) (see Listing 2). At run time, another user wanting to test an algorithm just provides the name of the problem to poli and gets an interface to a dynamically instantiated blackbox running in a different conda environment. The __call__ method then takes care of communication with the caller and logging to an observer. Some of these black boxes require additional assets (e.g. the weights of a neural network, or csv files). They are all either already provided or dynamically downloaded when the black box is used.
@@
Optimization algorithms can be oblivious to any requirements a black-box might have. Algorithms have only access to the AbstractBlackBox interface. They communicate with the problem only indirectly via a local network socket provided by python’s native multiprocessing library. This is the isolation mechanism that allows both algorithm and problem to run in different python environments.
Logging is done in general by observer objects. These are instantiated and attached to the problem, at which point they can log meta information on the problem and solver. Observers generally only observe black-box evaluations to track metrics and points which have been evaluated. We provide the option to also run an observer externally and in isolation, such that the environment is independent of problem and, or solver. Isolating the observer logic is essential to record comparable results across algorithms and problems, such that reported metrics are consistent. This also standardizes initial values for black-box observations.
We provide standard observers i.e. a standard mlflow observer and encourage the user to implement their own, should their needs require more complex logging logic for example reliant on metric computations.
A.5 Compute details for all experiments

### v2

@@ 4.2.2 Benchmarking HDBO on PMO
Hill | Genetic | CMAES | Hvarfner’s | Random | SAASBO | Turbo | Bounce | ProbRep | |
Oracle | Climbing | Algorithm | VanillaBO | LineBO | |||||
albuterol_similarity | |||||||||
amlodipine_mpo | |||||||||
celecoxib_rediscovery | |||||||||
deco_hop | |||||||||
drd2_docking | |||||||||
fexofenadine_mpo | |||||||||
gsk3_beta | |||||||||
isomer_c7h8n2o2 | |||||||||
isomer_c9h10n2o2pf2cl | |||||||||
jnk3 | |||||||||
median_1 | |||||||||
median_2 | |||||||||
mestranol_similarity | |||||||||
osimetrinib_mpo | |||||||||
perindopril_mpo | |||||||||
ranolazine_mpo | |||||||||
rdkit_logp | |||||||||
rdkit_qed | |||||||||
sa_tdc | |||||||||
scaffold_hop | |||||||||
sitagliptin_mpo | |||||||||
thiothixene_rediscovery | |||||||||
troglitazone_rediscovery | |||||||||
valsartan_smarts | |||||||||
zaleplon_mpo | |||||||||
@@
To benchmark the performance of HDBO on discrete sequences in continuous learned representations, we consider the PMO benchmark (Gao et al.,, 2022; Huang et al.,, 2021; Brown et al.,, 2019).666An introduction to PMO can be found in section A.5.2 We select Hvarfner’s VanillaBO, RandomLineBO, Turbo, BAxUS, SAASBO, Bounce, and ProbRep, including also HillClimbing, CMA-ES and GeneticAlgorithm as baselines. All solvers start from the same initial data to ensure a fair comparison except for BAxUS and Bounce due to the nature of their implementations. We test the aforementioned methods on PMO (Gao et al.,, 2022; Huang et al.,, 2021), which requires a discrete representation of small molecules. Thus, we train two MLP VAEs on SELFIES representations of small molecules using Zinc250k (Irwin et al.,, 2020; Zhu et al.,, 2022). These generative models had 2 and 128 latent dimensions, allowing us to get an impression of how these models scale with dimensionality. We restrict sequences to be of length 70 (adding [nop] tokens for padding); the post-processing renders an alphabet of 64 SELFIES tokens. Details can be found in Sec. A.3.
The average best result over 3 runs (of maximum 300 iterations each) is presented in Tables 2 and 3 for 128D and 2D latent spaces respectively. We see a clear advantage in the optimizers that work on learned representations, instead of in discrete space. Such a discrepancy is to be expected: methods that optimize in latent space have been presented with information prior to their optimization campaigns, while methods like Bounce and ProbRep explore the whole discrete space. Further, the simple baseline is reliably beaten by the continuous alternatives in lower dimensions except Turbo, but this advantage is not as clear in the 128D case, signaling a more complex problem. Some of these tasks, however, are equally challenging for all solvers. deco_hop remains close to the original default value of 0.5, and there is no improvement over valsartan_smarts (which only REINVENT improves on in the original PMO results (Gao et al.,, 2022; Loeffler et al.,, 2024)).
@@ 5 Conclusion
In this paper, we surveyed the field of high-dimensional Bayesian optimization (HDBO) focusing on discrete problems. This highlighted the need for (i) a novel taxonomy of the field that emphasizes the differences between methods that rely on unsupervised discrete information, and methods that optimize sequences directly, and (ii) a standardized framework for benchmarking HDBO methods. We approach these in the form of two software tools: poli and poli-baselines. Using these tools, we implemented several HDBO methods and tested them in a standard benchmark for sequence design and small molecule optimization. Our findings suggest that optimizers that work on discrete sequence space (i.e. the structured spaces family) perform as expected on problems of low complexity, but do not scale gracefully to problems with larger dictionaries/higher sequence lengths. In such cases, optimizers that leverage pre-trained latent-variable models have an edge over the other tested methods that work directly on sequence space. That being said, we find that, for low budgets, simpler baselines tend to perform as well or better than most BO methods, echoing recent criticisms of molecular optimization (Tripp and Hernández-Lobato,, 2023). Our framework opens the door to fair and easily-replicable comparisons. We expect poli-baselines to be used by practitioners for running HDBO solvers in up-to-date environments compared across several tasks in our ongoing benchmark, which we plan to expand to other discrete objectives in poli.
@@
Although we taxonomize different families of HDBO methods, we have only benchmarked a subset; moreover, the black boxes we use are often simple, closed-form functions or data-driven oracles, limiting the applicability of our benchmark to real-world scenarios. The field of discrete sequence optimization for biology/chemistry needs better black box oracles with more relevance to their respective domains. Another limitation of our approach is the lack of a systematic analysis on the myriad of design choices made on these optimization campaigns; we strived for documenting them properly, but an in-depth analysis is called for in future work. Our benchmark is ongoing and we plan to include further experiments in the project’s website with, hopefully, participation from the community. Finally, we note that optimizing small molecules or proteins opens the door to both drug discovery, but also dual use (Urbina et al.,, 2022).
@@
The work was partly funded by the Novo Nordisk Foundation through the Center for Basic Machine Learning Research in Life Science (NNF20OC0062606). RM is funded by the Danish Data Science Academy, which is funded by the Novo Nordisk Foundation (NNF21SA0069429) and VILLUM FONDEN (40516). SH was further supported by a research grant (42062) from VILLUM FONDEN as well as funding from the European Research Council (ERC) under the European Union’s Horizon programme (grant agreement 101125993). WB was supported by VILLUM FONDEN (40578). This work was in part supported by the Pioneer Centre for AI (DRNF grant number P1). MGD thanks Sergio Garrido and Anshuk Uppal for feedback on early versions of this document; Peter Mørch Groth for useful discussions; Luigi Nardi, Erik Hellsten and Raffaello Baluyot for feedback on the preprint version of this paper, and Samuel Stanton for feedback on the software.
@@ A.1 Methods Overview
closed-form ASM | September 22, 2020 | Wycoff et al., (2021) | ✓ |
@@
GaBO | November 22, 2021 | Jaquier et al., (2020) | ✓ |
@@
VAE-BO+Inv | July 22, 2022 | Verma et al., (2022) | ✗ |
Uncert | November 09, 2021 | Notin et al., (2021) | ✓ |
GAUCHE | September 21, 2023 | Griffiths et al., (2023) | ✓ |
LLMs | July 21, 2024 | Kristiadi et al., (2024) | ✓ |
AntBO | January 23, 2023 | Khan et al., (2023) | ✓ |
RMF | May 20, 2022 | Kim et al., (2022) | ✗ |
A.1.1 On how HDBO can be applied to discrete sequences
Fig. 4 shows an overview of the different ways in which HDBO could be applied to discrete sequences. On the left, we have the original sequence space with inputs being a list of categorical variables belonging to some alphabet; methods like Bounce, Prob. Rep and those that work on mixed/categorial inputs in the Structured Spaces family (see Sec. 3.7) work in this space. Since the original sequence space is large for the problems we are interested in, several other methods rely on subspaces from the original problem. Such subspaces can be constructed either as linear projections (Sec. 3.4) or as learned non-linear latent variables (Sec. 3.5). Once a latent continuous representation is available, all other families in the taxonomy become available and could be leveraged for latent space optimization. In these cases, there is usually a mapping between the subspace and the original high-dimensional space which allows for computing objective functions. In the linear case an up-projection is given by multiplying with a matrix of the right shape, and in the case of non-linear representations it is usually a decoding process.
In this paper, we explore applying the discrete optimizers of the Structured Spaces family in the sequence spaces of several Ehrlich problems (Sec. 4.2.1), as well as small molecule optimization using SELFIES representations. The continuous optimizers chosen from the other families are applied directly in one-hot space in the case of Ehrlich, or in the latent space of a Variational Autoencoder learned for SELFIES strings (Sec. 4.2.2).
@@ A.2 Reproducing results
We forked the official open source implementation of Bounce777https://github.com/LeoIV/bounce and added an interface between poli’s black boxes and their optimizer. Moreover, we made their implementation pip-installable. Bounce’s implementation is originally provided with an MIT license.
@@
provides an open-source implementation built on GPyTorch and BoTorch.888https://github.com/facebookresearch/bo_pr They provide a pip installable Python package which did not install until the dependencies mentioned above were fixed (to 1.11 and 0.7 respectively); further, the environment provided had to be updated by replacing the deprecated scikit-learn installation in a fork of their repository. After implementing an interface for poli black boxes, we relied on their script run_one_replication.py to implement a custom solver. The original PR code is provided with an MIT License.
@@
was implemented by following the tutorial on BoTorch,999https://github.com/pytorch/botorch/blob/main/tutorials/turbo_1.ipynb which can be found on GitHub under MIT License.
@@
is implemented using the Python package provided by the authors.101010https://github.com/LeoIV/BAxUS Since their work builds on the original TuRBO code from Uber, it inherits their license.
@@ A.3 Training VAEs on SELFIES
We use PyTorch to implement a VAE with [one_hot_input, 1024, 512, 256, latent_dim] encoder and a symmetric decoder, using ReLU activations. The latent space prior used is a standard Gaussian, and the decoded distribution is a categorical. We use torch.distributions to compute ELBO losses without any weighting.111111The exact implementation can be found here: https://github.com/MachineLearningLifeScience/hdbo_benchmark/blob/1a81f7fbd531eb4dc70f84828d6efc1f2ee53e5e/src/hdbo_benchmark/generative_models/vae_selfies.py. We are currently exploring a generative model with more expressive power, and plan to update the benchmarks accordingly in our project’s website.
@@
We train for a maximum of 1000 epochs using early stopping with a patience of 50 epochs. Our batch sizes are , learning rates are , and optimizer is AdamW.121212The training script can be found here: https://github.com/MachineLearningLifeScience/hdbo_benchmark/blob/master/src/hdbo_benchmark/experiments/training_vae_on_zinc250k/train_vae.py
@@ A.4 Technical details on poli and poli-baselines
This section gives a short overview on the possibilities of poli and poli-baselines. For further information please refer to the full online documentation under https://machinelearninglifescience.github.io/poli-docs/.
The aim of poli is to make it easier to benchmark new algorithms on new problems. There are three components to our framework: problems created by problem factories which contain black boxes, solvers that optimize these problems (aiming always at maximization), and observers which get attached to black boxes and log every single black box call. Fig. 5 shows these three components and their relationships. poli provides a unified numpy (Harris et al.,, 2020) interface, with inputs being numpy arrays of strings, and outputs being numpy arrays of floats. Finally, poli makes sure to log every call to the black box, as well as handling evaluation budgets—function calls are measured transparently and consistently across algorithms.
@@
A major issue on the endeavour of benchmarking on new problems are often conflicting package dependencies. When evaluating a brand-new solver on an older benchmark problem, it can simply be impossible to create a Python environment in which both can run. One may re-implement one or the other, though this is not only cumbersome but also error-prone. With the isolation mechanisms provided by poli this is no longer a problem. In particular, the isolation of the different components in the system (the black box, or the observer that is used for logging) allows an easy comparison to additional algorithms or on other problems, without the need to change plotting scripts.
LABEL:lst:example_experiment shows the general workflow:131313All listings in this appendix were tested using poli version v1.0.1, and poli-baselines version v1.0.2. They can be easily implemented and tested in a Colab notebook. a user instantiates a black box by calling create with the corresponding black box’s name, and the name of an observer. poli then takes care that both black box and observer are instantiated, in different conda environments if necessary. create returns an object inheriting from Problem which provides access to all things necessary: the black box function, initial observations, the observer and problem information like the alphabet and if sequences are aligned or not.141414A future release adds a data package attribute, attached to problems. The data package standardizes the information available to solvers by declaring unsupervised and supervised data; equalizing solver initialization. The black box is a function that takes and returns a numpy array, and the observer is informed automatically of any such calls. Each component of this workflow (black box, algorithm, and observer) can be easily exchanged by user-developed classes more specifically tailored to the need at hand. We refer to our online documentation151515https://machinelearninglifescience.github.io/poli-docs/ for the most recent information on how to do this.
To analyze an experiment a user has to define an observer, inheriting from AbstractObserver. LABEL:lst:example_observer shows how to quickly set up a simple text-file logger. The philosophy is to define the quantities of interest independent of problem or algorithm. We provide standard observers, for example an mlflow observer, and we encourage users to implement their own, to log other metrics of interest. Just as black boxes, observers can run in isolation, such that the environment is independent of problem and, or solver, to facilitate consistent recording of metrics. Whenever observers can run directly in the same environment as the solver, they could be attached directly to the black box using the set_observer method. LABEL:lst:example_observer_without_isolation shows an example using the same workflow as above, and the SimpleObserver presented in LABEL:lst:example_observer.
@@
To expand on the suite of available black boxes, the user has to implement a subclass of the core object of poli: AbstractBlackBox. The method to implement is _black_box, which takes as input an array of strings x (as well as an optional context), and outputs the result of evaluating it as an array of floats. Listing LABEL:lst:example_black_box shows an example code snippet. At run time, another user wanting to test an algorithm just imports the relevant black box object from poli’s repository, and gets an interface to a dynamically instantiated black box potentially running in a different conda environment. The __call__ method then takes care of communication with the caller and logging to an observer. Some of these black boxes require additional assets (e.g. the weights of a neural network, or csv files). They are all either already provided or dynamically downloaded when the black box is used.
@@
Optimization algorithms can be oblivious to any requirements a black box might have. Solvers have only access to the AbstractBlackBox interface. They communicate with the problem only indirectly via a local network socket provided by python’s native multiprocessing library. This is the isolation mechanism that allows both algorithm and problem to run in different python environments.
A.5 Benchmark Introduction
A.5.1 Ehrlich functions
Ehrlich functions (Stanton et al.,, 2024) were proposed as a closed-form alternative to black boxes like FoldX or RaSP, which could potentially raise licensing issues and, before poli, were not readily available for querying. This section explains how Ehrlich functions are constructed and queried from a bird’s eye view. The details can be found in the original paper.
These functions are procedurally generated from a random seed and hyperparameters like alphabet size , sequence length , number of motifs and motif length . The following are the steps that are followed to construct the problem:
1.
A sparse transition matrix of size is constructed. This transition matrix spans a finite Markov chain: sequences in the problem are constructed by starting with a random symbol and following the probabilities determined by . The rows of this transition matrix are probability vectors which determine how likely it is to go from one symbol in the alphabet to another. Such a transition matrix needs to satisfy that (i) the probability of going from one symbol to itself is non-zero, and (ii) it is always possible to go from any symbol to any other symbol after following a sequence of finite steps (irreducibility). In other words, we need the Markov chain spanned by to be ergodic. defines the search space of feasible sequence: the fact that is sparse means that some of the transitions are impossible.
2.
Once has been constructed, motifs of length are sampled according to it by sampling a sequence of length and splitting it. This ensures that the motifs are feasible and easily within reach from one to the next.
3.
These motifs are to be satisfied in certain positions in the string. These positions are constructed using random offsets of size . For example, the motif “ADER” with offsets of is satisfied if the sequence contains “ADXEXR” where “X” can be any other member of the alphabet: “D” is at distance 1 of “A”, “E” is at a distance 3 of “A” and “R” is at distance 5 of “A”. These random offsets are randomly constructed to maximize slackness in the sequence.
We take the original implementation161616https://github.com/prescient-design/holo-bench and wrap our black box logic around it, making it compatible with solvers in poli-baselines. We also include a data package made by using the original utilities for sampling the underlying transition matrix.
A.5.2 PMO
The Practical Molecular Optimization (PMO) benchmark contains a representative set of tasks defined on small molecule inputs with respective computational oracles. The input to the black box functions are alphabet representation for small molecules as either tokenized SMILES (Weininger,, 1988) or SELFIES (Krenn et al.,, 2020) (see below) for the exact alphabets used – in principle one can be converted into the other. While small molecule sequences usually contain fewer elements in a sequence than for example proteins, the alphabet can contain more tokens making this yet another high dimensional discrete optimization space. We build upon the work by Gao et al., (2022), who propose molecular optimization focused on validity, diversity, synthesizability - using computational values for all metrics and a budget of 10.000 evaluations. The PMO suite itself is based on the existing benchmarks Guacamol (Brown et al.,, 2019), and elements of the TDC (Huang et al.,, 2021). The types of optimization tasks for small molecules can be differentiated into: optimizing for simple metrics like QED, LogP171717These are aggregate properties of small molecules, and can sometimes be poor proxies for other chemical downstream behavior. qualitative tasks, such as scaffold hopping, rediscovery of particular substances (e.g. troglitazone), and more complex tasks such as docking surrogates and classifiers on fingerprint representations of molecules (GSK3, JNK3, DRD2) (Huang et al.,, 2021). We evaluate a selection of solvers across all tasks with multiple seeded runs on a fixed budget and report the average of the best observations, weighing all functions equally.181818Equal weighting can favor methods that perform well on a few exploitable tasks (e.g. logP, QED). We draw the users attention to that fact with the alternatives to discount or discard such tasks. Altogether, these tasks constitute a set of functions which can be optimized and which have previously been used to assess algorithm performance in the bio-chemical domain, allowing a comparisons to previous results and reported benchmarks. We make these environments accessible and solvable in a unified way through the poli infrastructure.
A.6 Compute details for all experiments
A.6.1 Ehrlich functions
The following solvers ran in an HPC with memory/CPU/GPU requirements given by 30G/10/1 Titan X with 12Gb of memory with: SAASBO, ProbRep, Hvarfner’s VanillaBO, RandomLineBO, DirectedEvolution, HillClimbing, CMA-ES and GeneticAlgorithm. All these were time-gated via SLURM to run for at most 24hrs.
The remaining solvers (BAxUS, Turbo and Bounce) ran in a Google Cloud instance with one Tesla T4 with 16Gb.
A.6.2 PMO